1. A 0.24 g sample of compound of oxygen and boron was found by analysis to contain 0.096 g of boron and 0.144 g of oxygen. Calculate the percentage composition of the compound by weight.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Total mass of compound = 0.24 g (Given) Mass of boron = 0.096 g (Given) Mass of oxygen = 0.144 g (Given) Thus, percentage of boron by weight in the 2 3 compound = 0.096/0.24 × 100% = 40% And, percentage of oxygen by weight in the = 0.144/0.24 × 100% = 60%